OpenStudy (anonymous):

MEDALS!!!!!!!! Systems of Equations A system of equations is shown below. 8x + 5y = 9 3x + 2y = 4 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

8x + 5y = 9 so you can eliminate variables, multiply this equation by -3 -24x + -15y = -27 3x+2y = 4 multiply this equation by 8 24x + 16y = 32 No combine the two equations: -24x + -15y = -27 24x + 16y = 32 y = 5 To find x, substitute 5 into the equations for y

OpenStudy (anonymous):

Would that be Part A?

OpenStudy (anonymous):

8x + 5(5) = 9 8x = -16 x =- 2 The solution to the system is (-2,5)

OpenStudy (anonymous):

So for example in Part B it would be 8(-2) + 5(5) = 9??

OpenStudy (anonymous):

That's part B

OpenStudy (anonymous):

How did you get that??

OpenStudy (anonymous):

It's ok, I was just wondering how you got those weird numbers

OpenStudy (anonymous):

Be sure to show that in both of the original equations

OpenStudy (anonymous):

Wait, is that Part B?? I thought you said that it was wrong??

OpenStudy (anonymous):

No..you were right. I was goofed up

OpenStudy (anonymous):

Ok thank you @DSS

OpenStudy (anonymous):

8(-2) + 5(5) = 9 -16 + 25 = 9 9 = 9 3x + 2y = 4 3(-2)+2(5)=4 -6+10=4 4 = 4 Part B